OpenStudy (linda):

3cos2theta-4cos^2theta+2=0. solve

OpenStudy (anonymous):

hey im pretty sure i got this one. you're just solving for x using identities, right?

OpenStudy (linda):

we r using substitutions to solve for diff. angle measures

OpenStudy (anonymous):

what do you mean by substitutions? i just used identites, which could be the same thing. i'm in trig now, so you might just be ahead of us. by i solved it...

OpenStudy (anonymous):

*but i solved it. do you want me to just go through what i did and see if thats what you're supposed to be doing?

OpenStudy (anonymous):

is it \[3\cos(2\theta)-4\cos^2(\theta) +2 =0 ?\]